What is the equation of the line described below written in slope-intercept form?
the line passing through point (-1, 5) and parallel to the line whose equation is x + y = 10

Answer by lwsshak3(11628)

What is the equation of the line described below written in slope-intercept form?
the line passing through point (-1, 5) and parallel to the line whose equation is x + y = 10
***
standard form of equation for a line in slope intercept form: y=mx+b, m=slope, b=y-intercept
given line
x+y=10
y=-x+10
slope, m=-1
..
slope of parallel line=-1
equation: y=-x+b
solve for b using coordinates (-1,5) of given point on the line.
5=-(-1)+b
b=4
equation: y=-x+4